OpenStudy (anonymous):

PLEASE HELP ME!!!! Let F(x)=2x+2,g(x)=x^2, and h(x)=3/x-2 State the domain of the operation. F(x)+g(x)

OpenStudy (amistre64):

Domain: (-inf,inf)

OpenStudy (amistre64):

what number makes this illegal to use? none.

OpenStudy (anonymous):

so -inf, inf is the answer?

OpenStudy (anonymous):

yeeeeeeeeeeeeeeeeeeeeeeees

OpenStudy (anonymous):

f(x) + g(x) = x^2 + 2x + 2. So the domain is \[(- \infty , \infty)\] But I don't see where the fn h(x) is used, does the question have more in it?